OpenStudy (anonymous):

If a is rational and b is irrational, is a + b necessarily irrational? What if a and b both are irrational?

OpenStudy (anonymous):

Lets see who can answer this one.

OpenStudy (nikita2):

1) b is irrational => b =p/q for some integers. a not = d/c for any integers. a + b = a + p/q = X. if X - rational, then b + X is rational too. So a+b irrational.

OpenStudy (nikita2):

2) Pi is irrational and - Pi is irrational too, so a+b may be rational.

OpenStudy (anonymous):

note that ur using wrong defintion, a number is a rational if it can be written as p/q where q is not equal to 0 and p and q are in their lowest terms

OpenStudy (anonymous):

the way u have proved it is absurd and not understandable at all

OpenStudy (anonymous):

Hint: prove part 1 by contradiction

OpenStudy (nikita2):

I see. Let's a is irrational, b is rational an X = a + b. If X is rational then X-b is rational too, but X-b=a contradiction

OpenStudy (anonymous):

i cant say if thats the right way for proving with contradiction

myininaya (myininaya):

Since a is rational, then a can be written as p/q,q not equal to 0 a+b=p/q+b=X p/q=b-X b-X can't be rational if X is rational since p/q is rational X has to be irrational i think i like nakita's form

OpenStudy (anonymous):

look at this, assume a + b is rational, then b = (a+b) - a which is rational since a is rational. But, b is irrational. A Contradiction. Hence the assumption is wrong. Therefore, a + b is irrational

myininaya (myininaya):

ok that is pretty lol

OpenStudy (nikita2):

very nice. I should learn this.

OpenStudy (anonymous):

nobody yet proved part 2